### v1

# Semantic Understanding and Data Imputation using Large Language Model to Accelerate Recommendation System

Abstract
This paper aims to address the challenge of sparse and missing data in recommendation systems, a significant hurdle in the age of big data. Traditional imputation methods struggle to capture complex relationships within the data. We propose a novel approach that fine-tune Large Language Model (LLM) and use it impute missing data for recommendation systems. LLM which is trained on vast amounts of text, is able to understand complex relationship among data and intelligently fill in missing information. This enriched data is then used by the recommendation system to generate more accurate and personalized suggestions, ultimately enhancing the user experience. We evaluate our LLM-based imputation method across various tasks within the recommendation system domain, including single classification, multi-classification, and regression compared to traditional data imputation methods. By demonstrating the superiority of LLM imputation over traditional methods, we establish its potential for improving recommendation system performance.
Index Terms:
Large Language Model, Data Imputation, Recommendation System
I Introduction
The exponential growth of big data has revolutionized many fields, offering unprecedented access to vast amounts of information. Researchers can find tons of information for uncovering patterns and making informed decisions [1, 2]. However, this abundance often masks a hidden adversary: sparse and small data. Missing information, often due to user inactivity, limited data collection, or technical constraints, can significantly hinder the effectiveness of big data models [3]. This is particularly true in recommendation systems, where personalized experiences hinge on a rich understanding of users and items, incomplete data significantly hinders the ability to generate accurate suggestions [4]. Traditional statistical methods for data imputation, like mean or median imputation, often fall short in capturing the complex relationships and underlying context within the data.
This paper tackles this challenge by proposing a novel approach that leverages the transformative power of Large Language Model (LLM) to address the challenge of data imputation in recommendation systems. LLMs, with their remarkable ability to process and understand vast amounts of natural language, possess the potential to intelligently fill in these missing data points. By harnessing the LLM’s capability to learn intricate relationships and context from large text corpora, our proposed method aims to impute data that is not only statistically sound but also semantically meaningful [5]. This enriched data can then be utilized by recommendation systems to generate more accurate and personalized suggestions for users.
Focusing on the domain of recommendation systems, we explore the specific application of LLM-based data imputation. Recommendation systems rely heavily on comprehensive user and item data to generate personalized suggestions that resonate with individual preferences. By effectively imputing missing values, we aim to create a more complete picture of user behavior and item characteristics. This, in turn, allows the recommendation system to generate more accurate and relevant suggestions, ultimately enhancing the user experience.
To comprehensively assess the effectiveness of our approach, we meticulously design a series of experiments that encompass a diverse range of classification and regression tasks. These experiments delve into single classification, where the system predicts a single category for an item, multi-classification, which allows for assigning multiple categories, and regression, where the focus is on predicting continuous values like ratings. By demonstrating the superiority of LLM imputation over traditional statistical methods across these varied scenarios, we aim to establish its significance as a game-changer in improving the performance of recommendation systems.
To comprehensively assess the effectiveness of LLM-based imputation, we conduct rigorous experiments across a diverse range of tasks within the recommendation system domain. These experiments encompass single classification, where the system predicts a single category for an item (e.g., AD recommendation), multi-classification, where multiple categories can be assigned (e.g., multiple categorical movies recommendation), and regression, which focuses on predicting continuous values like ratings or purchase likelihood (e.g., movie rating prediction). By demonstrating the advantage of LLM data imputation over traditional statistical methods in these varied scenarios, we experiment our proposed approach in different recommendations system tasks with different datasets. In summary, our paper makes the following primary contributions:
  • $\bullet$
    We propose a novel approach that utilize LLM to do data imputation which aims to handle sparse data and small data issue in big data model
  • $\bullet$
    We further utilize the imputed data and evaluate in the recommendation system which shows improvement to other statistical data imputation strategy
  • $\bullet$
    Extensive experiment are done to further verify LLM data imputation works better in single classification, multiple classification task and regression recommendation tasks.
II Related Work
II-A Data Imputation
Data imputation has been a well-established technique in statistics for handling missing data, with a rich history of methodological development. Early approaches focused on simple methods like mean or median imputation, which can introduce bias. More sophisticated techniques have since emerged, including k-Nearest Neighbors (kNN) imputation, which imputes missing values based on similar data points, and model-based methods that leverage statistical models to predict missing values [6]. Recent years, more researches are focus on machine learning algorithms for imputation, such as matrix factorization and deep learning techniques, which can handle complex patterns and relationships within the data for more accurate imputations. However, choosing the optimal imputation method depends on the characteristics of the data, the amount and pattern of missing data, and the intended analysis.
II-B Large Language Model
LLM is trained on massive amounts of text data, have shown promise due to their ability to capture complex relationships and semantic information within data. This capability allows them to potentially impute missing values in a more reliable way than traditional methods. For instance, some approaches treat imputation as a classification task, where the LLM predicts the most likely value for the missing entry based on the surrounding data. Others leverage the generative nature of LLMs to create a distribution of possible values, providing a more comprehensive picture of the imputation uncertainty. While promising, research on LLM-based data imputation is still evolving. Areas of exploration include mitigating potential biases present in training data and ensuring the imputed values maintain data integrity, particularly in sensitive domains like healthcare [7]. Overall, LLMs offer a new avenue for tackling missing data issues, with the potential to improve the accuracy and robustness of data analysis in various fields.
II-C Recommendation System
The goal of a recommendation system is to generate meaningful suggestions to a collection of users for items or products that might interest them. Data sparsity is a persistent challenge in such systems, significantly impacting the accuracy and effectiveness of recommendations. Collaborative filtering techniques, a mainstay in recommendation systems, struggle when user-item interaction matrices are highly sparse, with many missing entries. This sparsity makes it difficult to identify similar users or items for accurate recommendations [8]. More and more research has explored various approaches to address this issue, focus on developing robust recommendation systems that can effectively handle data sparsity and deliver personalized recommendations even with limited user-item interactions. In this paper, we aims to handle those missing data using LLM-based data imputation technology.
III Method
Figure 1: Framework of our proposed method. Some of users’ historical data are missing, fine-tune LLM can be utilized to impute the missing data. Complete data are used to feed into Recommendation System.
Figure 1 illustrates our proposed methodology. We leverage fine-tuning to adapt a large language model (LLM) on the complete data, excluding missing entries. Subsequently, we employ the fine-tuned LLM to impute missing values by providing the existing data as a prompt. A detailed explanation of this process will be provided in the following sections.
III-A Fine-tune LLM Model
In this work, we leverage the Low-Rank Adaptation (LoRA) technique  [9] to achieve efficient fine-tuning of LLM. LLM typically trained billions of parameters, rendering comprehensive fine-tuning computationally expensive. LoRA offers a cost-effective alternative by freezing the pre-trained model weights and introducing a set of trainable low-rank adapter parameters. This approach significantly reduces the computational burden associated with fine-tuning while enabling the LLM to adapt to the specific task or domain.
$W_{0}+\bigtriangleup W=W_{0}+BA$ (1)
where both A and B are trainable weights and $W_{0}$ is the freeze weights. In our case, we will use the data entries without any missing data to fine-tune the LLM model.
III-B Predict Missing Data
Subsequently, we employ the fine-tuned LLM for the task of missing data imputation. For each data instance containing one or more missing values, a prompt is constructed utilizing the existing data points to predict the missing information. For instance, given a data entry with attributes UserId=11, MovieId=22, Genres=NaN (indicating missing value), and Rating=4.5, the prompt would be formulated as: ”UserId is 11, MovieId is 22, Rating is 4.5, what is Genres?”. Additionally, the predicted value generated by the LLM will be used to replace the $NaN$ value.
III-C Evaluation in Recommendation System
To comprehensively assess the efficacy of the LLM-based data imputation approach, the newly constructed dataset was subsequently employed within a deep learning recommendation system. To achieve a holistic evaluation of the advantages offered by LLM-based imputation, performance metrics was utilized across various task categories, encompassing single classification, multi-class classification, and regression. Within the single classification domain, precision, recall, and F1-score were adopted as the evaluation metrics. For multi-class classification tasks, Recall at k (denoted as R@k) and Normalized Discounted Cumulative Gain at k (denoted as N@k) were employed. Finally, Mean Absolute Error (MAE), Mean Squared Error (MSE), or Root Mean Squared Error (RMSE) were leveraged to assess the performance of the regression task.
IV Experiment
IV-A Model and Dataset
We opted to leverage GPT-2 as our large language model (LLM) due to its open-source availability and demonstrably competent performance across diverse tasks. Regarding the datasets employed, we adopted a task-specific approach, tailoring the data distribution to the requirements of each individual task. We control the missing value rate of 5% enforced to mitigate potential biases and ensure model is suitable in different scenarios.
  • Single Classification: We leverage the AD Click dataset to evaluate the effectiveness of fine-tuned LLMs in imputing missing data. The imputed data is then fed into a recommendation system designed to classify user clicks on advertisements. This approach aims to improve the accuracy of predicting user engagement with targeted advertising.
  • Multiple Classification: We employ the widely used MovieLens dataset to assess the impact of LLM-based data imputation on movie recommendations. The imputed data is subsequently utilized by a recommendation system to suggest a personalized list of top-k movies for each user. This research aims to enhance the effectiveness of recommender systems by addressing data sparsity issues.
  • Regression: Building upon the MovieLens dataset, we investigate the use of LLMs for data imputation in predicting user ratings. The imputed data is then incorporated into a recommendation system tasked with predicting user ratings on a scale of 0.0 to 5.0. This approach seeks to improve the accuracy of rating predictions within recommender systems.
IV-B Data Imputation using Fine-tuned LLM
To fine-tune the LLM model, we pre-process the data and split the no-missing data and missing data entries. Data entries with no-missing data are used to fine-tune the LLM model. To impute missing values, we leverage the GReaT framework [10] to construct prompts for the LLM. These prompts incorporate the existing data within each row, allowing the LLM to predict the missing value based on the learned contextual relationships.
IV-C Evaluation
To better evaluate LLM-based data imputation technique, we evaluate our model’s performance across three tasks: single classification, multi-class classification, and regression. Two benchmark datasets, AD click and MovieLens, are employed. For each dataset, we meticulously curate the data to achieve a targeted missing value ratio of approximately 5%. Subsequently, we implement four data imputation strategies: traditional methods (mean, zero, KNN, and iterative) and our LLM-based approach. A baseline is established by using the data without imputation. Both imputed and non-imputed datasets are then fed into a recommendation system for training and evaluation. The DLRM [11] model is utilized for this purpose. Finally, a random split of 60/20/20 is implemented for training, testing, and validation, respectively.
IV-C1 Single Classification:
We assess the performance of various models for a single classification task using the AD Click dataset. We train a recommendation system model and evaluate its ability to predict user clicks on advertisements by comparing its predictions to ground truth labels. The detailed results are presented in Table I, with the top and second-highest performing models highlighted for clarity. While LLM-based data imputation approach did not achieve the absolute top performance in this particular task, as we will demonstrate in the following section, it exhibits potential for superior performance in more complex scenarios.
model $\uparrow$ $\uparrow$ f1$\uparrow$
baseline 0.1980 0.4450 0.2740
zero 0.7207 0.7200 0.7200
mean 0.8846 0.8700 0.8702
knn 0.9192 0.9150 0.9150
iterative 0.8970 0.8900 0.8903
LLM 0.9071 0.9001 0.9003
TABLE I: Comparison among LLM-based and statistical data imputation on single classification task
IV-C2 Multiple Classification:
The MovieLens dataset is used for evaluating model performance in a multiple classification task. Here, the objective is to recommend a user’s top $K$ movie preferences based on their historical data. Table II presents the results. Due to the richer metadata and intricate relationships within the MovieLens dataset, the LLM-based model demonstrates a clear advantage over other models.
model R@3$\uparrow$ R@5$\uparrow$ R@10$\uparrow$ N@3$\uparrow$ N@5$\uparrow$ N@10$\uparrow$
baseline 0.2510 0.3470 0.6050 0.4853 0.5381 0.7011
zero 0.2370 0.3250 0.5760 0.4412 0.5005 0.6629
mean 0.2610 0.3490 0.6110 0.5064 0.5455 0.7294
knn 0.2880 0.3870 0.6420 0.5213 0.5674 0.7331
iterative 0.2760 0.3680 0.6440 0.5154 0.5401 0.7420
LLM 0.2930 0.4050 0.6530 0.5692 0.6216 0.7632
TABLE II: Comparison among LLM-based and statistical data imputation on multiple classification task
IV-C3 Regression:
Finally, we evaluate the effectiveness of LLM-based data imputation within a regression task comparing with statistical methods. The MovieLens dataset is again utilized, but the goal here is to predict user ratings for individual movie-user pairs based on past interactions. Table III showcases the results, highlighting the superior performance of the LLM-based data imputation approach compared to other models.
[…]
baseline 0.7659 0.9792 0.9895
zero 0.7798 0.9928 0.9964
mean 0.7804 0.9883 0.9942
knn 0.7791 0.9909 0.9955
iterative 0.7785 0.9887 0.9943
LLM 0.7612 0.9647 0.9822
TABLE III: Comparison among LLM-based and statistical data imputation on regression task
V Conclusion
In conclusion, this paper proposes a novel approach that leverages the power of Large Language Models (LLMs) to address the challenge of data imputation in recommendation systems. By imputing missing data points in a semantically meaningful way, our method enriches data and allows recommendation systems to generate more accurate and personalized suggestions, ultimately enhancing user experience. We extensively evaluate our approach across various recommendation system tasks, demonstrating its effectiveness in improving performance compared to traditional data imputation methods. This research paves the way for utilizing LLMs to tackle sparse and small data issues in big data models, leading to more robust and informative recommendation systems.